A golf ball with a mass of 0.06 kg is subjected to an impulse that causes it to move with a velocity of 50 m/s. If the golf club and the ball are in contact for 5x10^-3 second, what average net force acts on the ball? 


1
Expert's answer
2022-04-06T13:59:41-0400

The Newton's second law says


"F=\frac{m\Delta v}{\Delta t}""F=\frac{0.06*50}{5*10^{-3}}=600\:\rm N"